Explain why in detail/using equations if possible!
media%2Ff8c%2Ff8cd7eed-bd61-49d4-8e23-f2
Answer #1

Solution) Bulbs in parallel glows brighter than bulbs in series as in parallel circuit same voltage flows through both the bulbs .

E = V(c) = V(D)

Where as in series voltage gets divided between the bulbs so here

V(A) = V(B) = E/2

as these are identical bulbs so same resistance and power is cobsidered .

So from above we can rank them as

Brightest to dimmest

C = D > A = B